The displacement x of a particle along a straight line at time t is given by x = a0 + a1 t+a2t2. The acceleration of the particle is

(a) a0

(b) a1

(c) 2a2

(d) a2

Grade:9

2 Answers

vikas askiitian expert
13 years ago

we know , v = dx/dt

             a = dv/dt

    x = ao + a1t + a2t2

v= dx/dt =  0 + a1 + 2a2t

a=dv/dt = d/dt (a1+2a2t)

            =0+2a2 

differntiation of constant is 0...

so accleration of particle = 2a2

option c is correct

Vamsi Krishna Gedela
13 years ago

x=a0+a1t +a2t2

accn=d2x/dt2=a2